two planes start from the same point and fly in opposite directions. the first plane is flying 50 mph slower than the second plane. in 2.5 hours the planes are 1400 miles apart. find the rate of each plane
Answer by mananth(16949)

Plane II x mph
Plane I x-50 mph
..
They are going away from each other.
So add up their speed.

combined speed = x+x-50
(2x-50)
Time = 2.5 hours
Distance = 1400 miles
Distance = speed * time
(2x-50)*2.5=1400
5x-125=1400
5x=1400+125
5x=1525
/5
x=305 mph Plane II
Plane I 305+-50 =255 mph